# DIP: Unsupervised Dense In-Context Post-training of Visual Representations

### v1

@@ 4.3 Ablations
In Tab. 6, we evaluate the impact of key design choices in our DIP approach. In-context vs Direct dense prediction.  In LABEL:tab:ablation_task, we compare our in-context dense prediction objective with direct dense prediction, where the model predicts pseudo semantic segmentation maps using a classification head. Our in-context approach significantly outperforms the direct method on PascalVOC and reduces RMSE on NYUv2 depth prediction, while both perform similarly on ADE20K.
Impact of “distractor” examples (LABEL:tab:ablation_distractor).
Removing the “distractor” support examples significantly reduces performance, as the absence of “distractors” simplifies the post-training task, limiting the model’s ability to learn discriminative dense representations.
Construction of positive examples.  In LABEL:tab:ablation_positives, we compare two strategies for creating positive examples: (1) retrieving nearest neighbors using DINOv2R image-wise features and (2) using two random crops from the same image. The nearest neighbor strategy outperforms random crops, validating our design choice and showing that our method avoids reliance on complex augmentations common in self-distillation and contrastive approaches.
Impact of DiffCut.  In LABEL:tab:ablation_pseudo_labels, we assess the role of DiffCut [15], which generates class-agnostic segments for pseudo segmentation maps before K-means clustering. Compared to applying K-means directly on DINOv2R dense features, DiffCut significantly improves performance, demonstrating its effectiveness as a training-free, unsupervised method leveraging Stable Diffusion. Furthermore, our pseudo labels achieve results nearly matching ground-truth semantic segmentation, underscoring their quality.
Impact of number of pseudo classes.  In LABEL:tab:ablation_num_pseudo_classes, we study the effect of the number of pseudo-classes (K-means clusters) used to generate pseudo-segmentation labels. Our approach shows robustness to this hyperparameter, with performance remaining stable across different values.
Post-training data: scene-centric vs object-centric.
In LABEL:tab:ablation_dataset, we compare DIP post-trained on the scene-centric COCO dataset versus the object-centric ImageNet dataset. Results are comparable, with COCO slightly outperforming ImageNet. This indicates that our method does not depend on human-curated, object-centric data and performs better with scene-centric data, which is easier to collect at scale.
@@ 5 Conclusion
We introduced DIP, a novel unsupervised post-training method that enhances dense image representations for in-context scene understanding. Leveraging meta-learning principles and automatically generating in-context tasks with pseudo-segmentations using Stable Diffusion, our approach avoids complex self-distillation architectures. DIP is computationally efficient ($<$9 hours on an A100 GPU) and generalizes well across downstream dense retrieval tasks, including semantic segmentation and depth prediction. It outperforms both the initial pretrained vision encoder and prior state-of-the-art post-training methods, providing a simpler and more effective solution for improving dense representations in vision models.
@@ Appendix A Implementation details
During post-training, we fine-tune only the last three transformer blocks of the pretrained ViT encoder $f(\cdot)$ while keeping the remaining layers frozen. Our MLP projector $h(\cdot)$ consists of two linear layers with a non-linear activation function GELU [35]. Hidden feature dimension is set to $7D$ and fixed output dimension 6144. $\ell_{2}$-normalization is applied at the output of the MLP. We set the temperature parameter $\tau$ of the softmax operator to $0.07$ (see Tab. 7). Our ablation study demonstrates that the method is robust to the choice of temperature, with values ranging from $0.03$ to $0.09$ yielding similar performance on the PascalVOC dataset. During post-training our support set consists of 1 positive and 7 “distractor” examples (8 total examples). During training, we use the AdamW [49] optimizer with a learning rate of $2.25\times 10^{-7}$ and a weight decay of $0.05$. We train for $5$ epochs. We employ cosine learning rate schedule.
Temperature 0.09 0.07 0.03
PascalVOC 80.8 81.0 81.0
@@ Appendix B Additional quantitative results
$10^{5}$$10^{6}$$10^{7}$3434343436363636383838384040404042424242Memory SizemIoU (%)(a) ViT-S/14 $10^{5}$$10^{6}$$10^{7}$3434343436363636383838384040404042424242Memory Size(a) ViT-B/14
@@ 1 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$
$1$ $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$ $1$ $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$

### v2

@@ 4.3 Ablations
In Tab. 6, we evaluate key design choices of DIP approach.
In-context vs Direct dense prediction (LABEL:tab:ablation_task).  We compare our in-context dense prediction objective with direct dense prediction, where the model predicts pseudo semantic segmentation maps using a classification head. Our in-context approach significantly outperforms the direct method on PascalVOC and reduces RMSE on NYUv2 depth prediction, while both perform similarly on ADE20K.
Impact of “distractor” examples (LABEL:tab:ablation_distractor).  Removing the “distractor” support examples significantly reduces performance, as the absence of “distractors” simplifies the post-training task, limiting the model’s ability to learn discriminative dense representations.
Construction of positive examples (LABEL:tab:ablation_positives).  We compare two strategies for creating positive examples: (1) retrieving nearest neighbors using DINOv2R image-wise features and (2) using two random crops from the same image. The nearest neighbor strategy outperforms random crops, validating our design choice and showing that our method avoids reliance on complex augmentations common in self-distillation and contrastive approaches.
Impact of DiffCut (LABEL:tab:ablation_pseudo_labels).  We assess the role of DiffCut [15], which generates class-agnostic segments for pseudo segmentation maps before K-means clustering. Compared to direct K-means on DINOv2R features, DiffCut significantly improves performance, demonstrating its effectiveness as a training-free, unsupervised method leveraging Stable Diffusion. Furthermore, our pseudo labels achieve results nearly matching ground-truth semantic segmentation, underscoring their quality.
Impact of number of pseudo classes (LABEL:tab:ablation_num_pseudo_classes).  We study the effect of pseudo-class count (K-means clusters) for pseudo-label generation. Our approach is robust to this hyperparameter, with stable performance across values.
Post-training data: scene-centric vs object-centric (LABEL:tab:ablation_dataset).  We compare DIP post-trained on scene-centric COCO vs. object-centric ImageNet. Results are comparable, COCO slightly outperforms ImageNet. This indicates that our method does not depend on human-curated, object-centric data and performs better on scene-centric data, easier to collect at scale.
@@ 5 Conclusion
We introduced DIP, a novel unsupervised post-training method that enhances dense image representations for in-context scene understanding. Leveraging meta-learning principles and automatically generating in-context tasks with pseudo-segmentations using Stable Diffusion, our approach avoids complex self-distillation architectures. DIP is computationally efficient ($<$9h on an A100) and generalizes well across downstream dense retrieval tasks, including semantic segmentation and depth prediction. It outperforms both the initial pretrained vision encoder and prior state-of-the-art post-training methods, providing a simpler and more effective solution for improving dense representations in vision models.
@@ Appendix A Implementation details
During post-training, we fine-tune only the last three transformer blocks of the pretrained ViT encoder $f(\cdot)$ while keeping the remaining layers frozen. Our MLP projector $h(\cdot)$ consists of two linear layers with a non-linear activation function GELU [35]. Hidden feature dimension is set to $7D$ and fixed output dimension 6144. $\ell_{2}$-normalization is applied at the output of the MLP. We set the temperature parameter $\tau$ of the softmax operator to 0.070.070.07 (see Tab. 7). Our ablation study demonstrates that the method is robust to the choice of temperature, with values ranging from 0.030.030.03 to 0.090.090.09 yielding similar performance on the PascalVOC dataset. During post-training our support set consists of 1 positive and 7 “distractor” examples (8 total examples). During training, we use the AdamW [49] optimizer with a learning rate of $2.25\times 10^{-7}$ and a weight decay of 0.050.050.05. We train for 555 epochs. We employ cosine learning rate schedule.
[…]
@@ Appendix B Additional quantitative results
$10^{5}$$10^{6}$$10^{7}$343434363636383838404040424242Memory SizemIoU (%)(a) ViT-S/14 $10^{5}$$10^{6}$$10^{7}$343434363636383838404040424242Memory Size(a) ViT-B/14
@@ 1 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$
111 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$ 111 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$